# Alignment via LLM-based Embedder

### v2

@@ Abstract
The original CLIP text encoder is limited by a maximum input length of 77 tokens, which hampers its ability to effectively process long texts and perform fine-grained semantic understanding. In addition, the CLIP text encoder lacks support for multilingual inputs. All these limitations significantly restrict its applicability across a broader range of tasks. Recent studies have attempted to replace the CLIP text encoder with an LLM-based embedder to enhance its ability in processing long texts, multilingual understanding, and fine-grained semantic comprehension. However, because the representation spaces of LLMs and the vision-language space of CLIP are pretrained independently without alignment priors, direct alignment using contrastive learning can disrupt the intrinsic vision-language alignment in the CLIP image encoder, leading to an underutilization of the knowledge acquired during pre-training. To address this challenge, we propose ProCLIP, a curriculum learning-based progressive vision-language alignment framework to effectively align the CLIP image encoder with an LLM-based embedder. Specifically, ProCLIP first distills knowledge from CLIP’s text encoder into the LLM-based embedder to leverage CLIP’s rich pretrained knowledge while establishing initial alignment between the LLM embedder and CLIP image encoder. Subsequently, ProCLIP further aligns the CLIP image encoder with the LLM-based embedder through image-text contrastive tuning, employing self-distillation regularization to avoid overfitting. To achieve a more effective alignment, instance semantic alignment loss and embedding structure alignment loss are employed during representation inheritance and contrastive tuning. Extensive experiments show ProCLIP achieves 6.8% to 13.5% improvement on zero-shot classification and presents excellent performance on cross-modal retrieval, multilingual cross-modal retrieval, and fine-grained understanding tasks, demonstrating the effectiveness and robustness of ProCLIP. The Code is available at https://github.com/VisionXLab/ProCLIP.
@@
CLIP demonstrates remarkable zero-shot recognition capabilities by learning joint vision-language representations through contrastive learning on large-scale image–text pairs (Radford et al., 2021). Serving as a bridge between vision and language, it is widely adopted in multiple downstream tasks such as image–text retrieval (Yang et al., 2023; Zheng et al., 2025), text-to-image generation (Wang et al., 2022), and open-vocabulary object detection (Wu et al., 2023b). However, the original CLIP model relies on English text captions with a maximum length of 77 tokens as its supervisory signal (Zhang et al., 2024). This design limits its capacity to process long-form text and restricts input to English-only (Tschannen et al., 2025). Additionally, due to the absence of supervision for fine-grained textual semantics, this limitation further impedes its semantic understanding capability (Hu et al., 2025; Gu et al., 2025a; b).
@@
To overcome these limitations, methods such as Long-CLIP (Zhang et al., 2024) interpolate positional embeddings and fine-tune on long text–image pairs to extend the input length. While effective for long-text understanding, such approaches still fall short in enhancing CLIP’s fine-grained semantic understanding and multilingual capabilities. Recently, LLM exhibits remarkable proficiency in natural language processing, and it has pivoted towards harnessing decoder-only architectures for effective representation learning (BehnamGhader et al., 2024; Lee et al., 2024). Following this trend, methods such as FLAME (Cao et al., 2025) and LLM2CLIP (Huang et al., 2024) propose to replace CLIP’s original text encoder with LLM-based embedders. By leveraging the rich open-world knowledge inherent in LLMs, these approaches aim to enhance CLIP’s representational capacity—particularly in processing longer and more complex image captions. However, these methods align the CLIP image encoder directly with the LLM-based text embedder through contrastive learning, while neglecting the rich pretrained knowledge within CLIP. This “from-scratch alignment” compels both encoders to learn a new representation space from scratch, disregarding the original CLIP alignment knowledge. Such an approach increases the risk of overfitting, particularly when training data is scarce, thereby compromising model generalization. This observation leads to a critical research question: How can we systematically leverage CLIP’s pretrained knowledge to achieve efficient cross-modal alignment with an LLM-based embedder while preserving generalization capability?
In this paper, we propose ProCLIP, a simple yet effective progressive vision-language alignment framework enhancing the CLIP. ProCLIP leverages curriculum learning to first guide the LLM-based embedder (only MLP trainable) to adapt to the CLIP text encoder’s representation space, and then uses contrastive learning to further learn joint image-text representations. Specifically, ProCLIP first distills knowledge from the original CLIP text encoder into the LLM-based embedder, establishing an initial alignment between the CLIP image encoder and LLM-based embedder. Subsequently, we conduct contrastive learning on image–text pairs to further improve this alignment. Since the LLM-based embedder is already partially aligned with the CLIP image encoder during the prior stage, the contrastive optimization process becomes more stable and preserves generalization more effectively. To further mitigate overfitting, we impose a self-distillation constraint on the CLIP image encoder throughout this stage, which stabilizes training and improves generalization. To prove the effectiveness of ProCLIP, we evaluate it on multiple tasks across diverse data scales and model sizes. Extensive experiment results demonstrate that ProCLIP achieves consistently significant improvements. The main contributions of this paper are summarized as follows:
  • We highlight the limitation of previous works: previous methods fail to fully exploit the pre-trained knowledge in CLIP, and their reliance on simplistic contrastive learning for cross-modal alignment significantly compromises CLIP’s inherent generalization capabilities.
  • We propose ProCLIP  a simple but effective Progressive vision-language alignment framework to enhance CLIP. ProCLIP initially distills the pretrained knowledge into the LLM-based embedder. After that, ProCLIP utilizes contrastive fine-tuning constrained by self-distillation to further enhance cross-modal alignment while preserving the model’s inherent knoweledge.
@@
  • We conduct extensive experiments on multiple tasks across diverse scales of data and model. Compared to the baseline, ProCLIP achieves 6.8% to 13.5% improvement on zero-shot classification and performs strongly on other tasks, including short-text cross-modal retrieval, long-text cross-modal retrieval, multilingual cross-modal retrieval, and fine-grained understanding.
Figure 1: Illustration of research gap. Previous work directly aligns the LLM-based embedder with the CLIP image encoder, disregarding the valuable knowledge embedded in the pre-trained CLIP model. In contrast, ProCLIP first transfers knowledge from CLIP’s text encoder to the LLM embedder via distillation, establishing an initial alignment. It then refines the alignment between the CLIP image encoder and the LLM-based embedder through image-text contrastive learning with self-distillation regularization.
@@
Vision-Language Contrastive Learning. Vision-language contrastive learning aims to learn robust multimodal representations by pretraining on large-scale image-text pairs. A seminal work in this area, CLIP (Radford et al., 2021) aligns visual and linguistic representations through contrastive learning, bridging both modalities in a shared semantic space. As a bridge between vision and language, CLIP has been widely applied in multimodal learning. It enables a variety of natural language-guided open-vocabulary recognition tasks, including image classification (Zhou et al., 2022b; a; Kim et al., 2024), open-vocabulary semantic segmentation (Ding et al., 2022; Li et al., 2022; Ghiasi et al., 2022; Xu et al., 2022; Cho et al., 2024; Lan et al., 2024), and open-vocabulary object detection (Du et al., 2022; Kaul et al., 2023). However, CLIP remains fundamentally constrained by its text encoder’s limited capacity and fixed input length, which hinders its ability to process multilingual and long texts and model fine-grained semantics. To mitigate these issues, several methods have been introduced. Long-CLIP (Zhang et al., 2024) extends the input length via positional embedding interpolation, yet still fails to capture nuanced semantic relationships. LoTLIP (Wu et al., 2024) incorporates corner tokens to aggregate diverse textual information, preserving short-text understanding while significantly improving performance on long texts. Nevertheless, constrained by the capabilities of the text encoder, LoTLIP cannot incorporate additional open-world knowledge and remains unable to handle multilingual inputs.
@@
LLMs for Representation Learning. Large language models have presented remarkable proficiency across a wide range of natural language processing tasks (Touvron et al., 2023; Achiam et al., 2023; Bai et al., 2023; Liu et al., 2024a). Recent research has pivoted towards harnessing decoder-only architectures for effective representation learning. For instance, LLM2Vec (BehnamGhader et al., 2024) converts pre-trained decoder-only LLMs into versatile text encoders by incorporating three principal advancements: bidirectional attention mechanisms, masked next-token prediction, and unsupervised contrastive alignment. Meanwhile, Qwen3-Embedding (Zhang et al., 2025b) capitalizes on the Qwen3 model’s strong multilingual understanding and generation abilities. By integrating a large-scale unsupervised pretraining and supervised fine-tuning on high-quality data, it achieves state-of-the-art performance on the MTEB benchmark (Muennighoff et al., 2022). Inspired by these advances, recent works (Huang et al., 2024; Cao et al., 2025; Zhang et al., 2025a) attempt to enhance CLIP by replacing its text encoder with a powerful LLM-based embedder, thereby improving its ability to process multilingual, longer, and more complex textual inputs. Although these approaches present promise, their alignment strategies remain overly coarse and often lead to degraded generalization. Developing more refined and effective alignment techniques thus remains a critical and open research challenge.
Knowledge Distillation. Knowledge distillation (Hinton et al., 2015) is widely used in deep learning to enhance model performance and reduce computational complexity. Typically, a larger teacher model transfers knowledge to a smaller student model by guiding the learning of features or output distributions. Alternatively, self-distillation methods enable knowledge transfer within a single model, where deeper layers supervise shallower ones (Zhang et al., 2019). In the context of CLIP, several distillation techniques have been introduced. TinyCLIP (Wu et al., 2023a) employs affinity mimicking to capture cross-modal interactions during distillation, allowing the student to replicate the teacher’s alignment behavior in a shared affinity space. CLIP-KD (Yang et al., 2024a) integrates multiple strategies—including relation-based, feature-based, gradient-based, and contrastive distillation—to maximize feature similarity between the teacher and student models. CLIP-CID (Yang et al., 2024b) utilizes cluster-instance discrimination to transfer semantic knowledge from the teacher, enabling the student to develop a richer understanding of the pretraining data. Different from the above method, this paper introduces a self-distillation mechanism to mitigate catastrophic forgetting during training and preserve the generalization capabilities of the model.
@@ 3.1 Preliminary
Contrastive Language-Image Pre-training. Contrastive Language-Image Pre-training (CLIP) (Radford et al., 2021) learns to align images and text from large-scale image–text pairs through contrastive learning, bridging both modalities in a shared embedding space. Given a batch of image-text pairs $\{(I_{i},T_{i})\}_{i=1}^{\mathcal{B}}$, the image encoder $\mathcal{E}_{I}$ and text encoder $\mathcal{E}_{T}$ map them into the joint semantic space as $\{(v_{i},t_{i})\}_{i=1}^{\mathcal{B}}$. To optimize both encoders in a dual-tower architecture, a symmetric contrastive learning objective is imposed on the resulting representations:
$\mathcal{L}_{CLIP}=-\sum_{i=1}^{\mathcal{B}}\Bigg[\underbrace{\log\frac{\exp(v_{i}\cdot t_{i}^{\top}/\tau)}{\sum_{j=1}^{\mathcal{B}}\exp(v_{i}\cdot t_{j}^{\top}/\tau)}}_{\text{text-to-image}}+\underbrace{\log\frac{\exp(t_{i}\cdot v_{i}^{\top}/\tau)}{\sum_{j=1}^{\mathcal{B}}\exp(t_{i}\cdot v_{j}^{\top}/\tau)}}_{\text{image-to-text}}\Bigg].$ (1)
@@
Improving CLIP with LLM-based Embedder. LLM2CLIP (Huang et al., 2024) first introduces an LLM-based embedder into CLIP, demonstrating enhanced long-text understanding. Given an LLM-based encoder $\mathcal{G}_{T}$, it encodes texts $\{T_{i}\}_{i=1}^{\mathcal{N}}$ offline into embeddings $\{t^{\prime}_{i}\}_{i=1}^{\mathcal{N}}$. This process is typically performed in an offline manner. During contrastive fine-tuning, a multilayer perceptron (MLP) is used to map $\{{t^{\prime}}\}_{i=1}^{\mathcal{N}}$ into the CLIP embedding space for dimensional alignment. The mapped text features and the image features from the CLIP image encoder are then optimized via the contrastive loss in Eq. 1, resulting in a newly aligned representation space. However, applying contrastive learning directly to fine-tuning data to optimize the MLP and vision encoder hinders the convergence of the new dual-tower architecture to an optimal parameter space. This arises because the text representations from the LLM-based embedder and MLP lack prior alignment with the vision encoder. Moreover, unconstrained fine-tuning may also cause excessive drift from the original pre-trained representation, while limited fine-tuning data (e.g., 3M samples) cannot compensate for the knowledge acquired during large-scale pre-training (e.g., 400M samples). To overcome these challenges, we propose a progressive alignment pipeline that improves multimodal alignment while preserving pre-trained knowledge.
3.2 Stage 1: Representation Inheritance via Cross-Architecture Distillation.
@@
Given a pre-trained image and text encoder of the CLIP model $\{\mathcal{E}_{I},\mathcal{E}_{T}\}$ and a pre-trained LLM-based embedder $\mathcal{G}_{T}$, our goal is to replace the CLIP text encoder $\mathcal{E}_{T}$ with the LLM-based embedder $\mathcal{G}_{T}$ to enhance comprehensive abilities. Consistent with prior works (Huang et al., 2024; Cao et al., 2025; Zhang et al., 2025a), we initially extract embeddings from textual captions offline using $\mathcal{G}_{T}$: $t^{\prime}=\{\mathcal{G}_{T}(T_{i})\in\mathbb{R}^{d}\}_{i=1}^{\mathcal{N}}$, where $d$ represents the embedding dimension of the LLM-based embedder.
@@
The embedding space of the LLM-based embedder exhibits no prior alignment with the CLIP image-text representation space. To bridge this gap, we adopt a cross-architecture distillation strategy that transfers knowledge from the CLIP text embedding space to the LLM embedding space. Specifically, given a batch of texts $\{T_{i}\}_{i=1}^{\mathcal{B}}$, we first utilize a MLP to unify the dimensions of LLM embeddings and CLIP text embeddings. To facilitate fine-grained semantic alignment, we propose an instance semantic alignment loss, denoted as $\mathcal{L}_{\text{ins}}$. This loss function leverages text-only data to distill knowledge from CLIP’s text encoder into the LLM-based embedder, defined as follows:
$\mathcal{L}_{\text{ins}}=\sum_{i=1}^{\mathcal{B}}\|\text{MLP}(t^{\prime}_{i})-\mathcal{E}(T_{i})\|_{2}.$ (2)
@@
Since $\mathcal{L}_{\text{ins}}$ only focuses on instance-level alignment without capturing the global embedding structure, we propose the embedding structure alignment loss $\mathcal{L}_{\text{struct}}$. This loss measures inter-sample distances within a batch in both the CLIP text encoder and LLM-based embedder spaces, and aligns the two globally by minimizing their pairwise distance discrepancy. $\mathcal{L}_{\text{struct}}$ is defined as:
$\mathcal{L}_{\text{struct}}=\sum_{\begin{subarray}{c}i,j=1\\ i<j\end{subarray}}^{\mathcal{B}}\big|\|\text{MLP}(t^{\prime}_{i})-\text{MLP}(t^{\prime}_{j})\|_{2}-\|\mathcal{E}(T_{i})-\mathcal{E}(T_{j})\|_{2}\big|.$ (3)
@@
3.3 Stage 2: Contrastive Tuning Integrated with Self-Distillation Regularization.
After the above phase, the $\text{MLP}(\mathcal{G}_{T})$ has already been preliminarily adapted to CLIP’s vision-language embedding space, making subsequent fine-tuning with vision-language contrastive learning significantly easier. We utilize the InfoNCE loss (Radford et al., 2021) to better align the image embedding $v_{i}$ and the projected LLM embedding $t_{i}^{*}=\text{MLP}(t^{\prime}_{i})$, which can be formulated as:
$\mathcal{L}_{\text{info}}=-\sum_{i=1}^{\mathcal{B}}\Bigg[\log\frac{\exp(v_{i}\cdot t_{i}^{*\top}/\tau)}{\sum_{j=1}^{\mathcal{B}}\exp(v_{i}\cdot t_{j}^{*\top}/\tau)}+\log\frac{\exp(t_{i}^{*}\cdot v_{i}^{\top}/\tau)}{\sum_{j=1}^{\mathcal{B}}\exp(t_{i}^{*}\cdot v_{j}^{\top}/\tau)}\Bigg],$ (4)
$\mathcal{L}_{\text{reg}}=\sum_{i=1}^{\mathcal{B}}\|\mathcal{E}_{I}(I_{i})-\mathcal{E}_{I}^{*}(I_{i})\|_{2}+\sum_{\begin{subarray}{c}i,j=1\\ i<j\end{subarray}}^{\mathcal{B}}\big|\|\mathcal{E}_{I}(I_{i})-\mathcal{E}_{I}(I_{j})\|_{2}-\|\mathcal{E}_{I}^{*}(I_{i})-\mathcal{E}_{I}^{*}(I_{j})\|_{2}\big|,$ (5)
where $\mathcal{E}_{I}^{*}$ denotes the EMA (Exponential Moving Average)-updated image encoder obtained as:
$\mathcal{E}_{I}^{*}=\alpha\mathcal{E}_{I}^{*}+(1-\alpha)\mathcal{E}_{I},$ (6)
where $\alpha$ controls the update rate of the teacher model parameters. The overall loss function of the contrastive tuning stage is defined as $\mathcal{L}_{\text{tune}}=\mathcal{L}_{\text{info}}+\lambda\mathcal{L}_{\text{reg}}$, where $\lambda$ is a loss weight.
4 Experiments
Method Data Flickr30k COCO ShareGPT4V Urban-1k DOCCI DCI Avg.
@@ 4 Experiments
ProCLIP 30M 94.5 81.6 69.3 53.2 96.8 97.0 93.0 94.4 87.5 89.8 75.9 79.5 86.2(+0.4) 82.6(+0.5)
Model Architecture: EVA02-CLIP ViT-L/14
EVA02-CLIP 2B 88.9 76.9 63.6 46.6 84.5 79.4 72.0 69.4 72.6 74.2 43.9 45.2 70.9 65.3
LLM2CLIP 3M 93.8 81.7 66.6 51.1 96.5 95.9 84.4 92.1 86.6 88.7 73.8 76.1 83.6 80.9
ProCLIP 3M 93.0 82.6 68.6 53.4 96.6 96.0 88.4 93.2 87.0 89.7 71.8 78.4 84.2(+0.6) 82.2(+1.3)
4.1 Experimental Setup
Datasets and Benchmarks. For the alignment dataset, we use CC3M (Changpinyo et al., 2021), CC12M (Changpinyo et al., 2021), and YFCC15M (Thomee et al., 2016),combined the high-quality captions from DreamLIP (Zheng et al., 2024). We conduct experiments with data scales of 3M (CC3M), 15M (CC3M + CC12M), and 30M (CC3M + CC12M + YFCC15M) to explore the effects of data scaling. For the benchmark, we perform zero-shot classification on 11 different classification datasets, robustness evaluations on 5 ImageNet variants, retrieval evaluations on 6 datasets, multilingual cross-modal retrieval evaluation on XM3600 (Thapliyal et al., 2022), and fine-grained understanding evaluation on MMVP-VLM (Tong et al., 2024). Regarding the model, we employ three OpenAI pre-trained CLIP models, ViT-B/32, ViT-B/16, and ViT-L/14, to investigate the effects of model scaling. Additionally, we conduct experiments with pretrained EVA02-CLIP (Fang et al., 2023) ViT-L/14 to assess the impacts of different model architectures. For the LLM-based embedder, we primarily use LLaMA3-8B-CC consistent with LLM2CLIP (Huang et al., 2024).
Implementation Details. For the representation inheritance phase, we train for four epochs, followed by another four epochs for contrastive tuning. During training, we employ AdamW (Loshchilov, 2019) as the optimizer, with a learning rate of $1\times 10^{-5}$ and a weight decay of 0.2. The parameters $\beta_{1}$ and $\beta_{2}$ are set to 0.9 and 0.98, respectively. In the first stage, the training batch size is set to 1024, while in the second stage it is increased to 4096. The loss weight $\lambda$ is set at 0.0004. Other training details can be found in the supplementary material.
4.2 Main Results
Cross-Modal Retrieval. As shown in Tab. 1, ProCLIP consistently surpasses LLM2CLIP in both short- and long-text retrieval tasks across various datasets and model scales. On short-text datasets such as Flickr30k and COCO, ProCLIP achieves significant improvements in both image-to-text (I2T) and text-to-image (T2I) retrieval. For instance, with ViT-L/14 and 30M training samples, it reaches 95.0% I2T Recall@1 on Flickr30k—nearly 2 percentage points higher than LLM2CLIP. On long-text benchmarks including DOCCI, DCI, and Urban-1k, ProCLIP also exhibits clear advantages. Under ViT-B/16 trained on 30M samples, it attains 73.6% (I2T) and 76.9% (T2I) on DCI. Moreover, across all data scales from 3M to 30M, ProCLIP delivers stable gains, with particularly strong improvements in T2I retrieval. These results confirm that ProCLIP enhances performance in both short- and long-text scenarios.

### v3

# ProCLIP: Progressive Vision-Language Alignment via LLM-based Embedder

@@ Abstract
Contrastive Language-Image Pre-training (CLIP) is constrained by its 77-token limit, lack of multilingual support, and coarse-grained semantic understanding. While replacing CLIP’s text encoder with a Large Language Model (LLM) offers a potential solution, direct contrastive alignment often disrupts the established cross-modal representation due to the lack of alignment priors between the two encoders. In this paper, we present ProCLIP, a curriculum learning framework designed to bridge LLMs and CLIP’s visual space, thereby unlocking CLIP’s potential for long-text, multilingual, and fine-grained understanding. The framework consists of two stages: ❶ Representation Inheritance, which distills CLIP’s original text-space knowledge into the LLM to establish an initial VL prior, and ❷ Contrastive Tuning, which refines the image-text alignment using self-distillation regularization to prevent catastrophic forgetting. To ensure semantic consistency, we introduce instance semantic alignment and embedding structure alignment losses throughout both stages. Extensive experiments demonstrate that ProCLIP improves zero-shot classification by 6.8%–13.5% and exhibits superior performance in long-text, multilingual, and fine-grained cross-modal retrieval tasks. The Code is available at https://github.com/VisionXLab/ProCLIP
ProCLIP: Progressive Vision-Language
Alignment via LLM-based Embedder
Xiaoxing Hu1,2*, Kaicheng Yang3*, Ziyang Gong1, Qi Ming4, Zonhao Guo5, Yu Tian 6, Xiang An3, Ziyong Feng3, Xue Yang1🖂 1Shanghai Jiao Tong University, 2Beijing Institute of Technology, 3DeepGlint 4Beijing University of Technology, 5Tsinghua University Equal contribution  🖂Corresponding author
@@ 1 Introduction
CLIP (Radford et al., 2021) establishes robust joint vision-language representations via large-scale contrastive learning, serving as a cornerstone for tasks like retrieval (Yang et al., 2023), generation (Wang et al., 2022), and detection (Wu et al., 2023b). Despite its success, CLIP is constrained by its 77-token English-only supervisory signal (Zhang et al., 2024; Tschannen et al., 2025). Furthermore, the lack of fine-grained textual supervision hampers its ability to capture nuanced semantic details (Hu et al., 2025).
Figure 1: Centered Kernel Alignment (CKA) analysis of visual representations. ProCLIP  outperforms the LLM2CLIP in retaining pretrained knowledge, providing a robust foundation for alignment.
@@
To overcome CLIP’s limitations, Long-CLIP (Zhang et al., 2024) interpolates positional embeddings for longer inputs, yet struggles with fine-grained and multilingual understanding. Meanwhile, FLAME (Cao et al., 2025) and LLM2CLIP (Huang et al., 2024) replace CLIP’s text encoder with LLM-based embedders (BehnamGhader et al., 2024; Lee et al., 2024) to leverage their rich linguistic priors. However, these methods typically perform "from-scratch alignment," directly forcing the image encoder to align with the LLM via contrastive learning. By neglecting CLIP’s original pretrained knowledge, this approach risks overfitting and undermines generalization, especially under data-scarce conditions. This motivates a critical question: How can we systematically leverage CLIP’s pretrained knowledge to achieve efficient cross-modal alignment with LLMs while preserving generalization?
We answer this question by proposing ProCLIP, a progressive alignment framework that reconceptualizes pretrained weights as foundational priors to bridge LLMs and CLIP’s visual space. ProCLIP adopts a two-stage curriculum:
❶ Representation Inheritance: We distill knowledge from the original CLIP text encoder into a lightweight adapter for the LLM. This anchors the frozen LLM to CLIP’s textual space, establishing a robust initial prior.
❷ Contrastive Tuning: Building on this foundation, we refine the joint representation space via contrastive tuning. A self-distillation constraint is applied to the image encoder to safeguard intrinsic vision-language knowledge and mitigate catastrophic forgetting.
We evaluated the preservation of pre-trained knowledge within the vision encoder following alignment, using Centered Kernel Alignment (CKA). As illustrated in Fig. 1, ProCLIP demonstrates superior retention of pre-trained features. This suggests that its carefully designed framework effectively facilitates cross-modal alignment without compromising generalization. ProCLIP  offers several advantages:1) Simplicity: An intuitive framework using pretrained weights as a semantic bridge. 2) Superiority: SOTA performance under identical data constraints. 3) Versatility: Extends CLIP to long-text, multilingual, and fine-grained scenarios with limited data. 4) Potential: A robust paradigm for modular upgrades in multi-encoder models.
Our main contributions are summarized as follows:
  • We identify a key limitation in current LLM-augmented CLIP models: "from-scratch alignment" neglects valuable pretrained priors, leading to compromised generalization.
  • We propose ProCLIP, a progressive framework that distills CLIP’s priors into an LLM-based embedder, followed by self-distilled contrastive tuning to refine alignment without catastrophic forgetting.
@@
  • We demonstrate ProCLIP’s efficacy through extensive experiments. ProCLIP achieves 6.8%–13.5% gains in zero-shot classification and shows significant improvements across short/long-text, multilingual retrieval, and fine-grained understanding tasks.
@@ 2 Related Work
Vision-Language Contrastive Learning. Vision-language contrastive learning, exemplified by CLIP (Radford et al., 2021), aims to align multimodal representations in a shared semantic space via large-scale pretraining. As a foundational bridge between modalities, CLIP enables diverse open-vocabulary tasks, including image classification (Zhou et al., 2022b, a; Kim et al., 2024), semantic segmentation (Ding et al., 2022; Li et al., 2022; Ghiasi et al., 2022; Xu et al., 2022; Cho et al., 2024; Lan et al., 2024), and object detection (Du et al., 2022; Kaul et al., 2023). Despite its versatility, CLIP is fundamentally constrained by its text encoder’s fixed input length and limited capacity, which hampers its performance in processing long-form, multilingual, and fine-grained semantics. To address these bottlenecks, Long-CLIP (Zhang et al., 2024) extends input lengths through positional embedding interpolation, while LoTLIP (Wu et al., 2024) aggregates diverse textual information using corner tokens to enhance long-text understanding. However, these methods remain tethered to the original CLIP text encoder’s architecture, preventing the integration of broader open-world knowledge and leaving multilingual support unresolved.
@@
LLM-based Representation Learning. Large Language Models (LLMs) have demonstrated exceptional proficiency across diverse natural language processing tasks (Touvron et al., 2023; Achiam et al., 2023; Bai et al., 2023; Liu et al., 2024a). Recent research has pivoted toward repurposing decoder-only architectures for robust representation learning. Notably, LLM2Vec (BehnamGhader et al., 2024) transforms LLMs into versatile encoders via bidirectional attention and contrastive alignment, while Qwen3-Embedding (Zhang et al., 2025b) leverages massive-scale pretraining to achieve state-of-the-art performance on the MTEB benchmark (Muennighoff et al., 2022). Inspired by these advancements, recent works (Huang et al., 2024; Cao et al., 2025; Zhang et al., 2025a) replace CLIP’s original text encoder with LLM-based embedders to handle multilingual, long-form, and complex textual inputs. However, these methods often employ coarse alignment strategies that inadvertently degrade the model’s inherent generalization. Addressing this gap by developing more refined and efficient alignment techniques remains a critical and open challenge.
Knowledge Distillation. Knowledge distillation (KD) (Hinton et al., 2015) facilitates knowledge transfer from a teacher to a student model, or within a single model via self-distillation (Zhang et al., 2019). In the vision-language domain, several KD techniques have been adapted for CLIP. TinyCLIP (Wu et al., 2023a) employs affinity mimicking to replicate cross-modal interactions, while CLIP-KD (Yang et al., 2024a) integrates feature-based and contrastive distillation to maximize teacher-student similarity. Furthermore, CLIP-CID (Yang et al., 2024b) utilizes cluster-instance discrimination to transfer semantic priors. Unlike these methods that primarily focus on model compression or performance enhancement, our work introduces a self-distillation mechanism specifically designed to mitigate catastrophic forgetting and safeguard the model’s intrinsic generalization during the transition to LLM-based architectures.
@@ 3.1 Preliminary
Contrastive Language-Image Pre-training. Contrastive Language-Image Pre-training (CLIP) Radford et al. (2021) learns to align images and text from large-scale image–text pairs through contrastive learning, bridging both modalities in a shared embedding space. Given a batch of image-text pairs $\{(I_{i},T_{i})\}_{i=1}^{\mathcal{B}}$, the image encoder $\mathcal{E}_{I}$ and text encoder $\mathcal{E}_{T}$ map them into the joint semantic space as $\{(v_{i},t_{i})\}_{i=1}^{\mathcal{B}}$. To optimize both encoders in a dual-tower architecture, a symmetric contrastive learning objective is imposed on the resulting representations:
$\begin{split}\mathcal{L}_{CLIP}=-\sum_{i=1}^{\mathcal{B}}\Bigg[&\underbrace{\log\frac{\exp(v_{i}\cdot t_{i}^{\top}/\tau)}{\sum_{j=1}^{\mathcal{B}}\exp(v_{i}\cdot t_{j}^{\top}/\tau)}}_{\text{text-to-image}}+\\ &\underbrace{\log\frac{\exp(t_{i}\cdot v_{i}^{\top}/\tau)}{\sum_{j=1}^{\mathcal{B}}\exp(t_{i}\cdot v_{j}^{\top}/\tau)}}_{\text{image-to-text}}\Bigg].\end{split}$ (1)
@@
Improving CLIP with LLM-based Embedder. LLM2CLIP Huang et al. (2024) first introduces an LLM-based embedder into CLIP, demonstrating enhanced long-text understanding. Given an LLM-based encoder $\mathcal{G}_{T}$, it encodes texts $\{T_{i}\}_{i=1}^{\mathcal{N}}$ offline into embeddings $\{t^{\prime}_{i}\}_{i=1}^{\mathcal{N}}$. This process is typically performed in an offline manner. During contrastive fine-tuning, a MLP-based adapter is used to map $\{{t^{\prime}}\}_{i=1}^{\mathcal{N}}$ into the CLIP embedding space for dimensional alignment. The mapped text features and the image features from the CLIP image encoder are then optimized via the contrastive loss in Eq. 1, resulting in a newly aligned representation space. However, applying contrastive learning directly to fine-tuning data to optimize the adapter and vision encoder hinders the convergence of the new dual-tower architecture to an optimal parameter space. This arises because the text representations from the LLM-based embedder and adapter lack prior alignment with the vision encoder. Moreover, unconstrained fine-tuning may also cause excessive drift from the original pre-trained representation, while limited fine-tuning data (e.g., 3M samples) cannot compensate for the knowledge acquired during large-scale pre-training (e.g., 400M samples). To overcome these challenges, we propose a progressive alignment pipeline that improves multimodal alignment while preserving pre-trained knowledge.
3.2 Stage I: Representation Inheritance via Cross-Architecture Distillation.
@@
Given a pre-trained image and text encoder of the CLIP model $\{\mathcal{E}_{I},\mathcal{E}_{T}\}$ and a pre-trained LLM-based embedder $\mathcal{G}_{T}$, our goal is to replace the CLIP text encoder $\mathcal{E}_{T}$ with the LLM-based embedder $\mathcal{G}_{T}$ to enhance comprehensive abilities. Consistent with prior works Huang et al. (2024); Cao et al. (2025); Zhang et al. (2025a), we initially extract embeddings from textual captions offline using $\mathcal{G}_{T}$: $t^{\prime}=\{\mathcal{G}_{T}(T_{i})\in\mathbb{R}^{d}\}_{i=1}^{\mathcal{N}}$, where $d$ represents the dimension of the LLM-based embedder.
@@
The embedding space of the LLM-based embedder exhibits no prior alignment with the CLIP image-text representation space. To bridge this gap, we adopt a cross-architecture distillation strategy that transfers knowledge from the CLIP text embedding space to the LLM embedding space. Specifically, given a batch of texts $\{T_{i}\}_{i=1}^{\mathcal{B}}$, we first utilize a single-layer MLP to unify the dimensions of LLM embeddings and CLIP text embeddings. To facilitate fine-grained semantic alignment, we propose an instance semantic alignment loss, denoted as $\mathcal{L}_{\text{ins}}$. Let $t_{i}^{*}=\text{Adapter}(t^{\prime}_{i})$ and $e_{i}=\mathcal{E}(T_{i})$ represent the projected LLM and original CLIP embeddings, respectively. This loss function leverages text-only data to distill knowledge from CLIP’s text encoder into the LLM-based embedder, defined as follows:
$\mathcal{L}_{\text{ins}}=\sum_{i=1}^{\mathcal{B}}\|t_{i}^{*}-e_{i}\|_{2}.$ (2)
@@
Since $\mathcal{L}_{\text{ins}}$ only focuses on instance-level alignment without capturing the global embedding structure, we propose the embedding structure alignment loss $\mathcal{L}_{\text{struct}}$. This loss measures inter-sample distances within a batch in both the CLIP text encoder and LLM-based embedder spaces, and aligns the two globally by minimizing their pairwise distance discrepancy. The structural loss is:
$\mathcal{L}_{\text{struct}}=\sum_{i<j}^{\mathcal{B}}\|d(t_{i}^{*},t_{j}^{*})-d(e_{i},e_{j})\|_{2},$ (3)
@@
3.3 Stage II: Contrastive Tuning Integrated with Self-Distillation Regularization.
After the above phase, the $\text{Adapter}(\mathcal{G}_{T})$ has already been preliminarily adapted to CLIP’s vision-language embedding space, making subsequent fine-tuning with vision-language contrastive learning significantly easier. We utilize the InfoNCE loss (Radford et al., 2021) to better align the image embedding $v_{i}$ and the projected LLM embedding $t_{i}^{*}$, which can be formulated as:
$\begin{split}\mathcal{L}_{\text{info}}=-\sum_{i=1}^{\mathcal{B}}\Bigg[&\log\frac{\exp(v_{i}\cdot t_{i}^{*\top}/\tau)}{\sum_{j=1}^{\mathcal{B}}\exp(v_{i}\cdot t_{j}^{*\top}/\tau)}+\\ &\log\frac{\exp(t_{i}^{*}\cdot v_{i}^{\top}/\tau)}{\sum_{j=1}^{\mathcal{B}}\exp(t_{i}^{*}\cdot v_{j}^{\top}/\tau)}\Bigg],\end{split}$ (4)
Method #Data Flickr30k COCO ShareGPT4V Urban-1k DOCCI DCI Avg.
@@
[…]